OpenStudy (anonymous):

Last week, Eva spent 8 and three-eighths hours on homework. This week, she spent ten and five-eighths hours. How many more hours did she spend this week?

OpenStudy (anonymous):

\[10\frac{5}{8} - 8\frac{3}{8} = 2\frac{2}{8}\]

OpenStudy (anonymous):

1 and three-fourths 2 and three-fourths 1 and one-fourth 2 and one-fourth these are the answers it gave me.

OpenStudy (anonymous):

2 and one-fourth